OpenStudy (anonymous):

1. Let f(x) = 49 – x2, g(x) = 7 – x. Find (f – g)(x).

OpenStudy (anonymous):

@amistre64 can you help me?

OpenStudy (anonymous):

f(x)-g(x)=49-x^2-(7-x)=49-x^2-7+x=-x^2+x+42

Parth (parthkohli):

\[(f - g)(x) \]is the same as\[f(x) - g(x) \]So, you have to find\[(49 - x^2)-( 7 - x) \]

Parth (parthkohli):

Are you sure it's not \((f \circ g)(x)\)?

OpenStudy (anonymous):

(f-g)(x) = f(x) - g(x), so substitute 49-x2 and 7-x into this equation. Try the rest yourself.